A highway study of 6,000 vehicles that passed by a checkpoint found that their speeds were normally distributed, with a mean of 110 Km/h and a standard deviation of 9 Km/h. Using the Empirical rule determine the percent of the vehicles had a speed between 83 and 92 km/h? The percent of vehicles with a speed between 83 and 92 km/h is: % (Omit the percent sign)
Answer #1

A highway study of 6000 vehicles that passes by a checkpoint found that their speed were normally distributed with mean of 110 km/hr & standard deviation of 9 km/hr.

So, \mu = 110, \sigma = 9

\mu - 3\sigma = 110 - 3(9) = 110 - 27 = 83, \mu + 3\sigma = 110 + 3(9) = 110 + 27 = 137

\mu - 2\sigma = 110 - 2(9) = 110 - 18 = 92, \mu + 2\sigma = 110 + 2(9) = 110 + 18 = 128

Now, By empirical rule -

95% of data will fall within two standard deviation from the mean, i.e. (\mu - 2\sigma, \mu + 2\sigma).

So, 95% of data will fall within (92,128).

This implies that about 5% of data fall below & above 92 & 128 respectively.

Since, normal distribution is symmetric distribution, hence 2.5% of the data fall below 92.

99.5% of data will fall within three standard deviation from the mean, i.e. (\mu - 3\sigma, \mu + 3\sigma).

So, 95% of data will fall within (81,137).

This implies that about 0.5% of data fall below & above 81 & 137 respectively.

Since, normal distribution is symmetric distribution, hence 0.25% of the data fall below 81.

Now, the percent of vehicles had speed between 83 & 92 km/hr -

P(83 < X < 92) = P(X<92) - P(X<83) = 2.5% - 0.25% = 2.25%

Hence, the percent of vehicles had speed between 83 & 92 km/hr is 2.25%.